# Highly Efficient Self-Adaptive Reward Shaping for Reinforcement Learning

### v1

@@ Abstract
Reward shaping addresses the challenge of sparse rewards in reinforcement learning by constructing denser and more informative reward signals. To achieve self-adaptive and highly efficient reward shaping, we propose a novel method that incorporates success rates derived from historical experiences into shaped rewards. Our approach utilizes success rates sampled from Beta distributions, which dynamically evolve from uncertain to reliable values as more data is collected. Initially, the self-adaptive success rates exhibit more randomness to encourage exploration. Over time, they become more certain to enhance exploitation, thus achieving a better balance between exploration and exploitation. We employ Kernel Density Estimation (KDE) combined with Random Fourier Features (RFF) to derive the Beta distributions, resulting in a computationally efficient implementation in high-dimensional continuous state spaces. This method provides a non-parametric and learning-free approach. The proposed method is evaluated on a wide range of continuous control tasks with sparse and delayed rewards, demonstrating significant improvements in sample efficiency and convergence stability compared to several baselines.
@@ 1 Introduction
Environments with sparse and delayed rewards present notable challenges for reinforcement learning (RL), despite its significant progress in various fields. In such contexts, as the reward structure lacks immediate signals, agents receive feedback only after long horizons, making the ability to quickly and effectively discover beneficial samples crucial for successful learning (Ladosz et al., 2022). To address this, a straightforward solution is to reconstruct the reward models by introducing auxiliary signals to assess the agent’s behavior at each step, which has led to the popular technique of Reward Shaping (RS) (Sorg et al., 2010a, b; Gupta et al., 2022). Inverse reinforcement learning (IRL), which extracts reward functions from human knowledge or expert demonstrations, represents an intuitive approach within this framework. However, this heavily relies on extensive human input and data, which can be difficult to obtain especially in complex environments. Alternatively, fully autonomous approaches have emerged as an attractive direction.
@@
Automatically maintained shaped rewards can be generally categorized into two branches: intrinsic motivation based signals, which are task-agnostic, and hidden value based signals, which are commonly task-related. The former mainly introduces exploration bonuses to encourage agents to explore a wider range of states, typically by rewarding novel or less frequently visited states (Ostrovski et al., 2017; Devidze et al., 2022; Tang et al., 2017; Burda et al., 2018; Bellemare et al., 2016). These methods have shown effective exploration enhancement, however, often overlooking the internal values of the states themselves. This can lead to the "noisy TV" problem, where agents become fixated on highly novel but meaningless regions, thus trapping agents in suboptimal behaviors (Mavor-Parker et al., 2022). In contrast, the latter leverages high-level heuristics to guide agents in retrieving hidden values from their experiences, which can significantly improve the agent’s convergence (Trott et al., 2019; Ma et al., 2024; Zheng et al., 2018).
@@
To overcome the limitations of existing RS methods, and try to combine the advantages of exploration-encouraged and hidden value-based rewards, this paper proposes a Self-Adaptive Success Rate based reward shaping mechanism (SASR) [^1] The source code is accessible at: https://github.com/mahaozhe/SASR. The success rate, defined as the ratio of a state’s presence in successful trajectories to its total occurrences, works as an auxiliary reward distilled from historical experience. The success rate assesses a state’s contribution toward successful task completion, which aligns with the agent’s original objectives, offering informative guidance for agent learning. Furthermore, to mitigate overconfidence caused by deterministic success rates, we adopt Beta distributions to model success rates from a probabilistic perspective. Beta distributions allow for a self-adaptive evolution of confidence in approximating a state’s success rate. This effectively prevents premature certainty and ensures stable convergence to reliable rewards as more data accumulates. To derive Beta distributions, we leverage kernel density estimation enhanced by random Fourier features to efficiently estimate success and failure counts, formulating a highly efficient approach. The main contributions of this paper are summarized as follows:
  1. (i)
@@
    We propose SASR, an autonomous reward-shaping mechanism without the requirement of expert demonstrations or domain knowledge. By extracting a metric from historical experiences that aligns with the agent’s optimization objectives, SASR effectively densifies the environmental rewards.
  2. (ii)
@@
    We introduce a novel self-adaptive strategy to achieve a balance between exploration and exploitation. Initially, low-confidence Beta distributions provide stochastic rewards that encourage exploration by diversifying the policy optimization directions. With the accumulation of experience, high-confidence Beta distributions offer more reliable and precise rewards to enhance exploitation.
  3. (iii)
@@
    To derive the Beta distributions in continuous state spaces, we implement kernel density estimation, leveraging random Fourier features. This formulates a non-parametric and learning-free approach, thereby achieving remarkably low computational complexity.
  4. (iv)
@@
    We use sparse-reward and continuous control environments, MuJoCo (Todorov et al., 2012) and simulated robot (de Lazcano et al., 2023), to evaluate the performance of SASR. It significantly outperforms several baselines in terms of sample efficiency, learning speed, and convergence stability.
@@ 2 Related Work
Reward shaping (RS) techniques can be broadly categorized based on the source of learning: either from human knowledge or agent’s own experiences. Techniques drawing reward models from human knowledge rely on pre-existing expertise, where methods such as Inverse Reinforcement Learning (IRL) (Arora and Doshi, 2021; Ramachandran and Amir, 2007; Ziebart et al., 2008; Hadfield-Menell et al., 2016) and Inverse Optimal Control (IOC) (Schultheis et al., 2021) aim to derive reward or objective functions from expert demonstrations. Another branch within this approach, Imitation Learning (IL) (Ho and Ermon, 2016; Duan et al., 2017; Brown et al., 2020), attempts to clone human behaviors from offline data without explicitly modeling reward structures. Following this, transferring the extracted reward models to new tasks has received considerable efforts (Bıyık et al., 2022; Wu et al., 2021; Ellis et al., 2021; Cheng et al., 2021; Adamczyk et al., 2023). However, these methods require significant human-driven data and generally show limited adaptability to out-of-distribution (OOD) scenarios. Therefore, we focus instead on the autonomous self-learning approaches, which based on the meaning of rewards, can be discussed from two perspectives: intrinsic motivation based and hidden value based rewards.
@@
Intrinsic motivation based RS explores general heuristics or task-agnostic metrics to encourage exploration in sparse-reward environments. Potential-based algorithms define the shaped reward in the form of $\gamma\Phi(s^{\prime})-\Phi(s)$, where $\Phi(\cdot)$ is a potential function, enabling the reward to be nullified in the Bellman equation, thus ensuring the optimal policy remains invariant (Devlin and Kudenko, 2012; Asmuth et al., 2008). However, the potential function’s design heavily depends on environmental dynamics, making it more suitable in model-based RL. More popular methods incorporate exploration bonuses by rewarding novel states (Devidze et al., 2022; Hong et al., 2018; Eysenbach et al., 2019). For instance, count-based strategies maintain visitation counts and assign higher rewards to less frequently visited states. To overcome the challenge of state counting in continuous spaces, Tang et al. (2017) utilized a hash function to discretize the state space, Bellemare et al. (2016) introduced pseudo-counts based on recording probabilities, and Ostrovski et al. (2017) used PixelCNN Van den Oord et al. (2016). Our work adopts a KDE approach to estimate the state counts. Additionally, Burda et al. (2018) proposed random network distillation to assess state novelty in high-dimensional spaces; Curiosity-driven approaches reward encounters with states that are surprising and out-of-prediction (Sun et al., 2022; Pathak et al., 2017). The exploration bonuses branch shows effective exploration enhancement, but only considering novelty while ignoring the underlying values may lead to sub-optimal behaviors. A notable example illustrating this limitation is the "noisy TV" problem (Mavor-Parker et al., 2022), where agents may become fixated on irrelevant and dynamic stimuli that yield high novelty but no useful information.
@@
Hidden value based RS, on the other hand, primarily focuses on task-related signals to reveal the extent to which states contribute to achieving higher environmental rewards and their underlying significance. Trott et al. (2019) integrated the distance between a state and the target as additional rewards. Stadie et al. (2020) derived informative reward structures using a Self-Tuning Network (STN) to optimize reward guidance. Memarian et al. (2021) captured the preferences among different trajectories by ranking them via a trained classifier. Zheng et al. (2018) minimized the KL-divergence between the learned and the original environmental rewards to synchronize their distributions. Mguni et al. (2023) used an auxiliary agent that competed against the original agent in a Markov game; Instead, Ma et al. (2024) introduced ReLara, a collaborative framework where an assistant reward agent automatically generates rewards to guide the policy agent. Moreover, the incorporation of multiple agents or hierarchical structures to share and transfer knowledge through synchronized reward functions is also a popular research direction (Park et al., 2023; Yi et al., 2022; Gupta et al., 2023; Hu et al., 2020).
@@
While intrinsic motivation rewards directly encourage exploration, hidden value rewards support efficient convergence, Ma et al. (2024) combined these two advances by introducing a scheme that evolves the shaped rewards from random perturbations to meaningful metrics. Injecting random rewards during early training relaxes the assumption that only novel states should be rewarded, suggesting instead that any state can be randomly rewarded. Building on this concept, our study considers the success rate, which is the probability that a state will contribute to task completion, as the hidden value. We sample this metric from evolving Beta distributions that progressively increase their confidence, thus aligning the process to balance the exploration and exploitation adaptively. Compared with ReLara, which relies on a black-box agent, our method derives the shaped rewards directly from historical data, effectively reducing the delay of reward convergence.
@@ 3 Preliminaries
Reinforcement Learning (RL) aims to train an agent to interact with an environment, which is commonly modeled as a Markov Decision Process (MDP). An MDP represented as $\langle S,A,T,R,\gamma\rangle$, involves four main components: $S$ is the state space, $A$ is the action space, $T:S\times A\times S\rightarrow[0,1]$ is the probability of transitioning from one state to another given a specific action, and $R:S\rightarrow\mathbb{R}$ is the reward model. The objective in RL is to learn a policy $\pi(a|s)$ that maximizes the expected cumulative rewards $G=\mathbb{E}[\sum_{t=0}^{\infty}\gamma^{t}R(s_{t})]$, where $\pi(a|s)$ indicates the likelihood of selecting action $a$ in state $s$, and $\gamma$ is the discount factor. Model-free RL methods include updating value functions, directly optimizing policies, or employing an actor-critic framework (Sutton and Barto, 2018).
@@ 4 Methodology
We propose a Self-Adaptive Success Rate based reward shaping mechanism (SASR) to accelerate RL algorithms in sparse-reward environments. Figure 1 illustrates the principles of the SASR mechanism: The diagram consists of two parts representing the early and late stages of learning. From the early stage to the late stage, the Beta distributions modeling the estimation of the true success rates evolve from being more stochastic to more deterministic as experiences accumulate. This progression closely aligns with the agent’s exploration-exploitation balance. The process of Beta distributions’ evolution and generation of shaped rewards from them are detailed in Section 4.1. Additionally, to achieve highly efficient computation, we leverage KDE and RFF to estimate success and failure counts, which are used to derive the corresponding Beta distributions, as detailed in Section 4.2. Lastly, Section 4.3 presents the integration of SASR into the RL agent and the overall algorithmic flow.
@@ 4.1 Self-Adaptive Success Rate Sampling
We consider the following augmented reward function for SASR:
$R^{SASR}(s)=R^{E}(s)+\lambda R^{S}(s),$ (3)
where $R^{E}(s)$ is the environmental reward, $R^{S}(s)$ is the additional shaped reward, and $\lambda$ is a weight factor to control its relative scale. When $\lambda\rightarrow 0$, we can recover the original reward model. We assign the shaped reward $R^{S}(s_{i})$ of a given state based on its success rate – defined as the ratio of the state’s presence in successful trajectories to its total occurrences. This metric provides a meaningful and informative reward from a statistical perspective, as a higher success rate indicates a greater likelihood and contribution for the state to guide the agent towards successful task completion, thus deserving a higher shaped reward. Formally, the success rate based shaped reward $R^{S}(s_{i})$ is given by:
@@
where $N_{S}(s_{i})$ and $N_{F}(s_{i})$ denote the counts of appearances for state $s_{i}$ in historical successful and failed trajectories respectively, and $f(\cdot)$ is a mapping function that adjusts the magnitude of rewards: $f:[0,1]\rightarrow[R_{\min},R_{\max}]$.
Given the counts $N_{S}(s_{i})$ and $N_{F}(s_{i})$ derived from prior experiences, directly using a deterministic success rate $\frac{N_{S}(s_{i})}{N_{S}(s_{i})+N_{F}(s_{i})}$ may lead to overconfidence in the estimation of the true value. To address this, we model the success rate estimation from a probabilistic perspective, inspired by the principles of Thompson sampling, an effective algorithm widely used to balance exploration and exploitation in multi-armed bandit problems (Thompson, 1933; Agrawal and Goyal, 2012). We model the success rate of each state as an approximated variable in a Beta distribution, setting the shape parameters $\alpha=N_{S}(s_{i})+1$ and $\beta=N_{F}(s_{i})+1$:
@@
where the beta function $B(\cdot,\cdot)$ is the normalization factor. By sampling from this distribution, we obtain a probabilistic estimate of the state’s success rate. This sampled value, $r^{S}_{i}$, is then processed through the mapping function $f(\cdot)$ to produce the shaped reward: $R^{S}(s_{i})=f(r^{S}_{i})$.
@@
  • Encourage Exploration. In the early stages of learning, the counts $N_{S}(s_{i})$ and $N_{F}(s_{i})$ are relatively low, resulting in broader Beta distributions and greater variability in sampled values, which injects significant randomness into the shaped rewards. These random rewards optimize the agent in various directions with tiny steps, thus shifting the anchors from which the stochastic policy samples actions. The strategy effectively broadens the range of policies and actions, thus encouraging the exploration. (Supporting experimental results are provided in Section 5.2.)
  • Enhance Exploitation. In the late stage of learning, as the counts $N_{S}(s_{i})$ and $N_{F}(s_{i})$ increase, the Beta distribution gradually sharpens, concentrating generated rewards closer to the true success rate. The evolution to a certain reward signal with higher confidence highly supports the agent’s exploitation phase, facilitating more efficient convergence towards optimal policies.
  • Consistent Optimization. The peak of the Beta distribution, $B(N_{S}(s_{i})+1,N_{F}(s_{i})+1)$, is given by $\frac{N_{S}(s_{i})}{N_{S}(s_{i})+N_{F}(s_{i})}$, directly approximating the success rate. Meanwhile, the expected value of the distribution, $\frac{N_{S}(s_{i})+1}{N_{S}(s_{i})+N_{F}(s_{i})+2}$, aligns closely with the calculated success rate. This ensures that, despite some randomness, the overall reward remains consistent with policy optimization.

### v3

@@ Abstract
Reward shaping is a technique in reinforcement learning that addresses the sparse-reward problem by providing more frequent and informative rewards. We introduce a self-adaptive and highly efficient reward shaping mechanism that incorporates success rates derived from historical experiences as shaped rewards. The success rates are sampled from Beta distributions, which dynamically evolve from uncertain to reliable values as data accumulates. Initially, the shaped rewards exhibit more randomness to encourage exploration, while over time, the increasing certainty enhances exploitation, naturally balancing exploration and exploitation. Our approach employs Kernel Density Estimation (KDE) combined with Random Fourier Features (RFF) to derive the Beta distributions, providing a computationally efficient, non-parametric, and learning-free solution for high-dimensional continuous state spaces. Our method is validated on various tasks with extremely sparse rewards, demonstrating notable improvements in sample efficiency and convergence stability over relevant baselines.
@@ 1 Introduction
Environments with extremely sparse rewards present notable challenges for reinforcement learning (RL). In such contexts, as the reward model lacks immediate signals, agents receive feedback only after long horizons, making the ability to quickly discover beneficial samples crucial for successful learning (Ladosz et al.,, 2022). To address this, a straightforward solution is to reconstruct the reward models by introducing auxiliary signals to assess the agent’s behavior, which has led to the popular technique of Reward Shaping (RS) (Strehl and Littman,, 2008; Gupta et al.,, 2022). Inverse reinforcement learning (IRL), which extracts reward functions from human knowledge or expert demonstrations, represents an intuitive approach within this framework Arora and Doshi, (2021). However, IRL heavily relies on extensive human input, which can be difficult to obtain especially in complex environments. Alternatively, fully autonomous approaches have emerged as an attractive direction.
@@
Automatically maintained reward shaping can be broadly categorized into two branches: intrinsic motivation-based rewards, which are task-agnostic, and inherent value-based rewards, which are typically task-specific. The former mainly introduces exploration bonuses to encourage agents to explore a wider range of states, commonly by rewarding novel or infrequently visited states (Burda et al.,, 2018; Ostrovski et al.,, 2017; Tang et al.,, 2017; Bellemare et al.,, 2016). While these methods effectively enhance exploration, they tend to overlook the internal values of the states. This can lead to the "noisy TV" problem, where agents fixate on highly novel but meaningless regions, thus trapping them in suboptimal behaviors (Mavor-Parker et al.,, 2022). In contrast, the latter leverages high-level heuristics to guide agents in extracting meaningful values from learning experiences, which helps stabilize convergence. (Ma et al., 2024b, ; Memarian et al.,, 2021; Trott et al.,, 2019).
@@
To overcome the limitations of existing RS methods and combine the advantages of exploration-encouraged and inherent value-based rewards, this paper introduces a novel Self-Adaptive Success Rate based reward shaping mechanism (SASR). The success rate, defined as the ratio of a state’s presence in successful trajectories to its total occurrences, works as an auxiliary reward distilled from historical experience. This success rate assesses a state’s contribution toward successful task completion, which closely aligns with the agent’s original objectives, offering informative guidance for learning. Furthermore, to mitigate overconfidence caused by deterministic success rates, we adopt Beta distributions to model success rates from a probabilistic perspective. Beta distributions enable a self-adaptive evolution of confidence in approximating a state’s success rate, ensuring the system gradually converges to reliable rewards as more data is collected, while preventing premature certainty. To derive Beta distributions, we use kernel density estimation (KDE) enhanced by random Fourier features (RFF) to estimate success and failure counts, formulating a highly efficient approach. The main contributions of this paper are summarized as follows:
@@
  • We propose SASR, an autonomous reward-shaping mechanism without the requirement of domain knowledge. By extracting a metric from historical experiences that aligns with the agent’s optimization objectives, SASR effectively augments the environmental rewards.
@@
  • A novel self-adaptive mechanism is introduced to achieve an efficient exploration-exploitation balance. Initially, low-confidence Beta distributions provide stochastic rewards that encourage exploration by diversifying policy optimization. As more experience accumulates, high-confidence Beta distributions offer more reliable and precise rewards to enhance exploitation.
@@
  • To derive the Beta distributions in continuous state spaces, we implement KDE with RFF. This formulates a non-parametric and learning-free approach, thereby achieving remarkably low computational complexity.
@@
  • SASR is evaluated on various extremely sparse-reward and continuous control tasks, significantly outperforming several baselines in sample efficiency, learning speed, and convergence stability.
@@ 2 Related Work
Reward shaping (RS) methods can generally be categorized based on the source of learning: either from human knowledge or agent’s own experiences. Techniques that derive reward models from human knowledge, such as Inverse Reinforcement Learning (IRL) (Arora and Doshi,, 2021; Ramachandran and Amir,, 2007; Ziebart et al.,, 2008; Hadfield-Menell et al.,, 2016) and Inverse Optimal Control (IOC) (Schultheis et al.,, 2021), aim to extract reward or objective functions from expert demonstrations. Following this, transferring the learned reward models to new tasks has received considerable efforts (Bıyık et al.,, 2022; Wu et al.,, 2021; Ellis et al.,, 2021; Cheng et al.,, 2021; Adamczyk et al.,, 2023). However, these methods rely heavily on human-generated data and often struggle to adapt to out-of-distribution scenarios. Thus, our focus shifts toward autonomous self-learning approaches, which can be further divided into intrinsic motivation-based and inherent value-based rewards based on the nature of the rewards.
@@
Intrinsic motivation based RS explores general heuristics or task-agnostic metrics to encourage exploration. Potential-based algorithms define the shaped reward as $\gamma\Phi(s^{\prime})-\Phi(s)$, where $\Phi(\cdot)$ is a potential function. This ensures the reward cancels out in the Bellman equation, preserving the optimal policy (Devlin and Kudenko,, 2012; Asmuth et al.,, 2008; Wiewiora,, 2003). However, designing the potential function is highly dependent on the environmental dynamics, making it more applicable to model-based RL. More commonly, methods incorporate exploration bonuses to reward novel states (Devidze et al.,, 2022; Badia et al.,, 2020; Hong et al.,, 2018; Eysenbach et al.,, 2019). Representatively, count-based strategies track visitation counts and assign higher rewards to less frequently visited states (Lobel et al.,, 2023; Machado et al.,, 2020; Fox et al.,, 2018; Fu et al.,, 2017; Martin et al.,, 2017). In continuous spaces, state counting is challenging, so Tang et al., (2017) introduced a hash function to discretize the state space, Bellemare et al., (2016) proposed pseudo-counts based on recording probabilities, and Ostrovski et al., (2017) used PixelCNN Van den Oord et al., (2016). Additionally, random network distillation based methods measure state novelty by neural networks (Yang et al.,, 2024; Burda et al.,, 2018), curiosity-driven approaches reward agents for encountering surprising or unpredictable states (Sun et al.,, 2022; Burda et al.,, 2019; Pathak et al.,, 2017). Although intrinsic motivation has proven effective in enhancing exploration, only considering novelty while ignoring the inherent values of states may lead to suboptimal policies. A notable example illustrating this limitation is the "noisy TV" problem (Mavor-Parker et al.,, 2022), where agents may become attracted to irrelevant, dynamic stimuli that generate high novelty but provide no useful information.
@@
Inherent value based RS, on the other hand, focuses on task-related signals that highlight how states contribute to achieving higher rewards and their underlying significance. For instance, Trott et al., (2019) introduced additional rewards based on the distance between a state and the target; Stadie et al., (2020) derived informative reward structures using a Self-Tuning Network to optimize guidance; Memarian et al., (2021) captured the preferences among different trajectories by ranking them via a trained classifier; Zheng et al., (2018) minimized the KL-divergence between learned and original rewards to align their distributions. Mguni et al., (2023) used an auxiliary agent competing against the original agent in a Markov game; Instead, Ma et al., 2024b introduced ReLara, a collaborative framework where an assistant reward agent automatically generates rewards to guide the policy agent. Moreover, incorporating multiple agents or hierarchical structures to share and transfer knowledge through synchronized reward functions is another promising research direction (Gupta et al.,, 2023; Ma et al., 2024c, ; Ma et al., 2024a, ; Park et al.,, 2023; Ma et al.,, 2023; Yi et al.,, 2022; Hu et al.,, 2020; Raileanu and Rocktäschel,, 2020).
@@
While intrinsic motivation rewards guide exploration, inherent value rewards support convergence, Ma et al., 2024b combined these two advances by introducing a scheme that evolves the shaped rewards from random perturbations into meaningful metrics. Injecting random rewards during early training relaxes the assumption that only novel states should be rewarded, suggesting instead that any state can be randomly rewarded. Building on this idea, our study considers the success rate as an informative value. We sample this metric from evolving Beta distributions that progressively increase in confidence, enabling a self-adaptive exploration-exploitation balance. Compared with ReLara, which relies on a black-box agent, our method draws information from historical data, effectively reducing the delay of reward convergence and enhancing explainability.
@@ 3 Preliminaries
Reinforcement Learning (RL) aims to train an agent to interact with an environment, which is commonly modeled as a Markov Decision Process (MDP). An MDP represented as $\langle S,A,T,R,\gamma\rangle$, involves four main components: $S$ is the state space, $A$ is the action space, $T:S\times A\times S\rightarrow[0,1]$ is the probability of transitioning from one state to another given a specific action, and $R:S\rightarrow\mathbb{R}$ is the reward model. The objective in RL is to learn a policy $\pi(a|s)$ that maximizes the expected cumulative rewards $G=\mathbb{E}[\sum_{t=0}^{\infty}\gamma^{t}R(s_{t})]$, where $\pi(a|s)$ indicates the likelihood of selecting action $a$ in state $s$, and $\gamma$ is the discount factor (Sutton and Barto,, 2018).
@@ 4 Methodology
We propose a Self-Adaptive Success Rate based reward shaping mechanism (SASR) to accelerate RL algorithms learning in extremely-sparse-reward environments. Figure 1 illustrates the principles of the SASR mechanism: The diagram consists of two parts representing the early and late learning stages. As experiences accumulate with learning progresses, the Beta distributions modeling the success rates evolve from being more stochastic to more deterministic. This autonomous adaption closely aligns with the agent’s exploration-exploitation balance. Section 4.1 introduces how Beta distributions evolve and how shaped rewards are generated from them. Additionally, to achieve highly efficient computation, we leverage KDE and RFF to estimate success and failure counts, which are used to derive the corresponding Beta distributions, as detailed in Section 4.2. Lastly, Section 4.3 presents the integration of SASR into the RL agent and the overall algorithmic flow.
@@ 4.1 Self-Adaptive Success Rate Sampling
We formulate the augmented reward function in SASR as adding an auxiliary shaped reward $R^{S}(s)$ to the environmental reward $R^{E}(s)$, weighting by a factor $\lambda$:
$R^{SASR}(s)=R^{E}(s)+\lambda R^{S}(s).$ (3)
We assign the shaped reward $R^{S}(s_{i})$ of a given state based on its success rate – defined as the ratio of the state’s presence in successful trajectories to its total occurrences. This metric provides a meaningful reward from a statistical perspective: a higher success rate, reflected in a higher shaped reward, indicates a greater likelihood that the state will guide the agent toward successful task completion. Formally, the success rate based shaped reward $R^{S}(s_{i})$ is given by:
@@
where $N_{S}(s_{i})$ and $N_{F}(s_{i})$ denote the counts of state $s_{i}$ appearing in successful and failed historical trajectories, respectively, and $f(\cdot)$ is a mapping function that adjusts the magnitude of rewards: $f:[0,1]\rightarrow[R_{\min},R_{\max}]$.
Given $N_{S}(s_{i})$ and $N_{F}(s_{i})$, directly using a deterministic success rate $N_{S}(s_{i})/(N_{S}(s_{i})+N_{F}(s_{i}))$ may lead to overconfidence in the estimation of the true value. To address this, we model the success rate estimation from a probabilistic perspective, inspired by the principles of Thompson sampling, an effective algorithm widely used to balance exploration and exploitation in multi-armed bandit problems (Thompson,, 1933; Agrawal and Goyal,, 2012). We model the success rate of each state as an approximated variable in a Beta distribution, setting the shape parameters $\alpha=N_{S}(s_{i})+1$ and $\beta=N_{F}(s_{i})+1$:
@@
where the beta function $B(\cdot,\cdot)$ is the normalization factor. By sampling from this distribution, we obtain a probabilistic estimate of the true success rate. This sampled value, $r^{S}_{i}$, is then processed through the scaling function $f(\cdot)$ to produce the shaped reward: $R^{S}(s_{i})=f(r^{S}_{i})$.
@@
  • Encourage Exploration. In the early learning phases, the counts $N_{S}(s_{i})$ and $N_{F}(s_{i})$ are relatively low, resulting in broader Beta distributions and greater variability in sampled values, which injects significant randomness into the shaped rewards. These random rewards optimize the agent in various directions with tiny steps, thus shifting the anchors from which the stochastic policy samples actions. The strategy effectively broadens the range of policies and actions, thus encouraging the exploration. (Supporting experimental results in Section 5.2.)
  • Enhance Exploitation. In the late learning phases, as the counts $N_{S}(s_{i})$ and $N_{F}(s_{i})$ increase, the Beta distribution gradually sharpens, concentrating generated rewards closer to the true success rate. The evolution to a certain reward signal with higher confidence highly supports the agent’s exploitation, facilitating more efficient convergence towards optimal policies.
  • Consistent Optimization. The peak of the Beta distribution, computed as $N_{S}(s_{i})/(N_{S}(s_{i})+N_{F}(s_{i}))$, directly equals to the success rate. Meanwhile, the expectation, $(N_{S}(s_{i})+1)/(N_{S}(s_{i})+N_{F}(s_{i})+2)$, closely approximates the success rate. This ensures that, despite stochastic, the overall reward remains consistent with policy optimization.